OpenStudy (anonymous):

Given the function T(z) = z – 6, find T(–4).

OpenStudy (anonymous):

to find T(-4) substitute -4 for z in the equation. Let me know if you need more help.

OpenStudy (anonymous):

can you show me, i dont understand

OpenStudy (anonymous):

T(z) = z-6 T(-4) = (-4)-6 T(-4) = -4-6 T(-4) = -10

OpenStudy (anonymous):

ohhh i get it now! thanks alot!

OpenStudy (anonymous):

you r welcome